OpenStudy (anonymous):

Heather hits a tennis ball upward with an initial speed of 32 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

h = 0 when the ball hit the ground, so 0 = rt - 16t^2 Since r = 32, 0= 32t - 16t^2 16t(2 - t) = 0 t = 0 or 2 At t=0, h=0 but the ball hasn't left the ground yet. Thus, the answer is 2. I hope this can help.. :))

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

welcome.. :)))

OpenStudy (anonymous):

And thanks for the medal..